# Systematics of neutron emission

### v2

@@ Abstract
Neutron physics is one of the oldest branches of the experimental nuclear physics, but the investigation of the spontaneous neutron emission from the ground state along the neutron dripline is still at its beginning, in spite of the crucial importance for nuclear astrophysics. The proton dripline is much better investigated and a systematics of spontaneous proton half lives is available in terms of an exponential dependence upon the Coulomb parameter $\chi\sim Q^{-1/2}$, characterizing the outgoing irregular Coulomb wave in terms of the Q-value (Geiger-Nuttall law). Our purpose is to propose a similar simple systematics of spontaneous neutron half lives, but in terms of the nuclear reduced radius $\rho\sim Q^{1/2}$, characterizing the ”neutral” outgoing irregular Neumann wave. We evidence the important role of the angular momentum carried by the emitted neutron. The influence of the neutron wave function generated by a Woods-Saxon nuclear mean field is also analyzed.
@@ 1 Introduction
In Ref. [10] a systematics of spontaneous proton emission half lives in terms of the Coulomb parameter ($\chi$) was proposed by defining the outgoing Coulomb-Hankel wave
@@
with a given angular momentum $l$. This function practically coincides with the irregular Coulomb wave $G_{l}(\chi,\rho)$ inside the barrier. It turns out that this function squared is proportional to the proton half life [1]. For neutron emission the Coulomb parameter vanishes, but the other parameter, defining the outgoing ”neutral” irregular Neumann function, is given by the reduced radius $\rho$. The aim of this letter is to propose a similar systematics of spontaneous neutron half lives, but in terms of this quantity.
@@ 2 Theoretical background
The theory of neutron emission is an old issue, but we will remind its main details in order to fix the main concepts. The standard way to analyze a single particle (proton or neutron) unbound state is to consider it as a resonant state in continuum. This is an eigenstate of some mean field potential with complex energy and outgoing asymptotics [1]. For outgoing narrow resonances, called Gamow states, the real part of the energy is positive and it is called Q-value. The decay width $\Gamma$ is twice the complex part of the energy and for narrow resonances it satisfies the condition $\Gamma<<Q$. Thus, the wave function describing the emission process is given by the following ansatz
@@
defining the S-matrix $S_{l}(E)=\exp[2i\delta_{l}(E)]$, the phase shift for some narrow resonant state sharply passes through the value $\delta_{l}=\pi/2$ where only the irregular component $G_{l}$ remains. Thus, for a narrow resonance both complex Gamow and real scattering state descriptions give close results.
The low-lying proton eigenstates in continuum have very small decay widths and therefore are long living due to the large Coulomb potential. In this case the equivalent expression of the particle decay width can be expressed by using the continuity equation as follows [1]
$\Gamma_{lj}=\hbar v\frac{s_{lj}}{\Omega_{j}}|N_{lj}|^{2}~{},$ (2.3)
where $l,j$ denotes the angular momentum and total spin of the emitted particle, $s_{lj}$ the spectroscopic factor, $\Omega_{j}=j+\frac{1}{2}$ the level occupancy and $v=\sqrt{2Q/\mu}$ the particle velocity with the reduced mass $\mu$. For superfluid nuclei the spectroscopic factor is given by the probability of the particle component in the quasiparticle state at the Fermi level $s_{lj}=u^{2}_{lj}\approx$ 0.5. The quantity $N_{lj}$ is called scattering amplitude and it is given by the ratio
$N_{lj}=\frac{f_{lj}(R)}{G_{l}(\chi,\rho)/\rho}~{},$ (2.4)
between the internal wave function normalized to unity inside nucleus and external outgoing wave, given by the irregular Coulomb wave, depending upon the Coulomb-Sommerfeld parameter $\chi=2Ze^{2}/(\hbar v)$, where Z is the daughter charge, and nuclear reduced radius $\rho=\kappa R$ defined in terms of the linear momentum $\kappa=\mu v/\hbar$. Let us mention that the above ratio obviously does not depend upon the radius $R$ in the region where the nuclear interaction vanishes, because both internal and external functions satify the same Schrödinger equation. Let us mention that the same result is obtained within the real scattering approach [28].
At the same time, some low-lying neutron resonant states have relative large centrifugal barriers and therefore are also relative long living. Therefore the decay width can be expressed by (2.3), but the external outgoing wave for chargeless particles with $\chi=0$ is given by the spherical Neumann function $G_{l}(0,\rho)/\rho\equiv y_{l}(\rho)$ and describes the neutron motion at large distances, where the nuclear interaction vanishes.
Figure 1: The values of $|C_{l}(\rho)|^{-2}$, where the centrifugal factor $C_{l}$ is given by Eq. (2.6), for $l$=1, 2, 3.
The spherical Neumann together with spherical Bessel function define the spherical Hankel function
@@
It turn out that it can be factorized into monopole and centrifugal components as follows
$h^{(\pm)}_{l}(\rho)$ $=$ $\frac{e^{\pm i\rho}}{\rho}~{}~{}\sum_{m=0}^{l}\frac{i^{m-l}}{m!(2\rho)^{m}}\frac{(l+m)!}{(l-m)!}$ (2.6)
$\equiv$ $h^{(\pm)}_{0}(\rho)C_{l}(\rho)~{}.$
Let us mention that monopole components with $C_{0}=1$ have the following simple expressions
$j_{0}(\rho)=\frac{\sin\rho}{\rho}~{},~{}~{}~{}y_{0}(\rho)=\frac{\cos\rho}{\rho}~{}.$ (2.7)
In Fig. 1 we plotted the inverse of the centrifugal factor squared $|C_{l}(\rho)|^{-2}$ for $l$=1 (solid line), $l$=2 (dashed line) and $l$=3 (dot-dashed solid line) versus the reduced radius $\rho$ estimated at the geometrical nuclear radius $R_{0}=1.2A^{1/3}$
$\rho=\kappa R_{0}=\frac{(2mQ)^{1/2}}{\hbar}1.2A^{1/3}\approx 0.263~{}Q^{1/2}A^{1/3}~{},$ (2.8)
written in terms of the neutron Q-value and mass-number $A$. Notice that by considering $Q<$ 2 MeV, $A<$ 30 one obtains a relative small interval of values $\rho<$ 1.
@@ 3 Data systematics
At present the amount of neutron emission experimental data from the ground state is very limited, due to the difficulties to measure half lives of these very unstable nuclei, close to the neutron dripline [29]. Moreover, the experimental errors are relative large. This is seen from Table I, where we give the available experimental data concerning quantum numbers, Q-value, half-life and error of neutron emission from ground state. In the last columns are given the corresponding references. Notice that, except for the first three lines, we assigned angular momenta according to the standard spherical shell model scheme [30], due to the fact that this quantity in general is not experimentally determined. Let us mention here that the standard closure numbers N=2, 8, 20 seems to change with the increase of the asymmetry $N-Z$. As can be seen from Fig. 1 of Ref. [31], the neutron shell closure at N=20 disapears and therefore the assignement of the angular momentum may be different from the standard scheme. Thus, we considered one exception labeled by an asterisk in Table I for $l=1$, corresponding to a much better fit to the theoretical curve, describing the angular momentum dependence of the half life.
TABLE I. Neutron experimental emission data: charge, neutron, mass numbers, angular momentum, Q-value, reduced radius estimated at the nuclear geometrical radius, half-life, error and reference. $Z$ $N$ $A$ $l$ $Q$ (keV) $\rho$ $T$ (s) $\delta T$ (s) Ref. 1 3 4 1 1600 0.528 8.42(-23) - [33] 2 3 5 1 735 0.386 7.89(-22) - [34] 2 5 7 1 409 0.322 3.04(-21) 0.41(-21) [34] 4 9 13 2 510 0.442 2.70(-21) 1.80(-21) [35] 4 11 15 2 1800 0.870 7.87(-22) 2.71(-22) [36] 5 11 16 1 82 0.190 $>$4.56(-21) - [37] 5 15 20 2 1560 0.892 9.12(-22) - [38] 8 17 25 2 725 0.655 2.28(-22) ${}^{+6.84(-22)}_{-2.28(-22)}$ [39] 9 19 28 2 220 0.375 4.56(-20) - [40] 1 219 0.373 2.53(-21) ${}^{+0.72(-21)}_{-0.46(-21)}$ [41]
@@
These nuclei belong to the neutron dripline. They are very unstable, the largest half life being only by two orders of magnitude above the characteristic nuclear time $10^{-22}$ s. Let us stress on the fact that the amount of experimental data is very rich if one considers neutron emission from excited states populated by beta decay, defining the so-called beta delayed neutron emission [32]. Let us mention here the main difference with respect to usual emission from single particle neutron resonances in the nuclear mean field. The delayed emission is a more complex phenomenon, namely it takes places from $1^{+}$ Gamow-Teller proton-neutron collective states in odd-odd nuclei, or their coupling to the isobaric analog state (IAS) in odd mass nuclei.
@@
Figure 2: Experimental half lives in logarithmic scale for $l$=1 (open circles) and $l=2$ (dark circles). The fitting curves are given by $T_{l}=0.1~{}y^{2}_{l}(\rho)$ for angular momenta l=1,2. We also added the curve corresponding to $l=0$.
@@
In Fig. 2 we plotted the experimental half lives together with the simplest theoretical prediction given by $T=\hbar\ln{2}/\Gamma$, where $\Gamma$ is given by (2.3), fitting these data by using one common parameter for all angular momenta
$T_{l}=0.1~{}y^{2}_{l}(\rho)~{}.$ (3.1)
Thus, we neglected the influence of the internal wave function $f_{lj}(R)$ and level occupancy $\Omega_{j}=j+\frac{1}{2}$ in Eq. (2.3). We determined the coefficient by using the experimental data with known angular momentum $l=1$. We notice a rather good agreement between this prediction and experimental data within the experimental errors.
@@
Figure 3: Experimental half lives in logarithmic scale (in $10^{-22}$ units) divided by the centrifugal factor $|C_{l}(\rho)|^{2}$. The fitting curve represents the same ansatz as in Fig. 2 for the monopole case $T_{0}=0.1~{}y^{2}_{0}(\rho)$.
@@
Let us mention that in Ref. [10] the influence of the centrifugal barrier for proton emitters was excluded by dividing the experimental half life to the centrifugal factor $C^{2}_{l}(\rho)$. Here we will use the same centrifugal factor by defining the reduced half life
$T_{red}=\frac{T_{exp}}{C^{2}_{l}}~{}.$ (3.2)
@@
It actually correspons to a dependence of the reduced half life upon the monopole spherical Neumann function squared. In Fig. 3 we plotted the dependence of the reduced half life in logarithmic scale versus versus the reduced radius $\rho$, together with the fitting curve $0.1~{}y^{2}_{0}(\rho)$. We mention here the quasilinear dependence of the curve versus $\rho$ over the experimental range. We obtained a rather good correlation between experimental points, except two points corresponding to the largest reduced radius.
@@
Figure 4: Experimental formation probability given by the ratio $T_{exp}/[0.1y_{l}^{2}(\rho)]$ in logarithmic scale.
@@
Let us here mention that the role of the particle emission formation probability was extensivelly investigated. In the recent review [13] various versions of the shell model calculations in this region were analyzed, starting with the Kuo-Brown code [42]. The particle emission processes were analyzed by using the resonant states in a deformed Woods-Saxon plus Coulomb mean field [43, 44, 45], as well as more sophisticated open quantum system formalism for the description of weakly bound nuclei far from the valley of stability [46, 47].
@@
In order to evidence the role of the neutron wave function we plotted in Fig. 4 the so-called ”experimental formation probability”, given by the ratio between the experimental half life and the best fit of the centrifugal penetration factor $f_{exp}(\rho)=T_{exp}/[0.1y_{l}^{2}(\rho)]$. This quantity is obviously proportional to the neutron wave function squared $f^{2}_{lj}$ entering the definition of the decay width (2.3)-(2.4). The result in Fig. 4 is nothing else that the ratio between experimental points and the corresponding values of the curve in Fig. 3. Notice that the rms error is $\sigma=0.492$, corresponding to a factor about three in the half life. Therefore the role of the neutron wave function is not so important for a rough estimate of the neutron emission half life for transitions connecting ground states.
@@
Figure 5: Logarithm of the half life versus energy of l=2 (upper curve) and l=1 (lower curve) resonant levels in a spherical Woods-Saxion mean field for 28F. The vertical line represnts the experimental Q-value and the two horizontal lines the experfimental half lives given by the last two lines of Table I.
@@
On the other hand, it turns out that the most sensitive ingredient is the angular momentum of the emitted neutron, similar to the proton emission systematics [10]. This is shown in Fig. 2 and also in Fig. 5, where we plotted the theoretical half life of a resonant state above the Fermi surface with angular momentum l=1 (lower curve) and l=2 (lower curve) versus its energy. Notice here two orders of magnitude difference in half lives between these curves at the experimental Q-value. These states were generated by using a spherical Woods-Saxon meand field [1] with universal parametrisation [48, 49] for 28F. We used the theoretical decay width given by (2.3) and changed the energy of the resonant state by using the real part of the central potential. It turns out that this relation is not sensitive to the change of other parameters of the used nuclear mean field. This feature is also general for proton emission processes connecting ground states [11]. By vertical dotted line we denoted the experimental Q-value and by horizontal dotted lines the two experimental values of the half life mentioned in the last two lines of Table I. Let us mention that our theoretical prediction for l=1 is much closer to the experintal value, thus confirming the above statement aboout the island of inversion in this region and magicity of 28O.

### v3

@@ Abstract
Neutron physics is one of the oldest branches of the experimental nuclear physics, but the investigation of the spontaneous neutron emission from the ground state along the neutron dripline is still at its beginning, in spite of the crucial importance for nuclear astrophysics. The proton dripline is much better investigated and a systematics of spontaneous proton half lives corected by the centrifugal barrier (monopole transitions) is given by the Geiger-Nuttall law $\log_{10}T\sim\chi$, where $\chi\sim ZQ^{-1/2}$ is the Coulomb parameter characterizing the outgoing Coulomb-Hankel wave in terms of the daughter charge $Z$ and Q-value. Our purpose is to propose a similar simple systematics of spontaneous neutron half lives, but in terms of the nuclear reduced radius $\rho=\kappa R\sim A^{1/3}Q^{1/2}$, characterizing the ”neutral” outgoing spherical Hankel wave. It turns out that the half life in emission of neutral particles is governed by the scaling law $T\sim\rho^{-2}\sim A^{-2/3}Q^{-1}$ for monopole transitions. We evidence the important role of the angular momentum carried by the emitted neutron. The influence of the neutron wave function generated by a Woods-Saxon nuclear mean field is also analyzed.
@@ 1 Introduction
In Ref. [10] a systematics of spontaneous proton emission half lives was proposed in terms of the Coulomb parameter ($\chi$), defining the outgoing Coulomb-Hankel wave
@@
depending on irregular $G_{l}(\chi,\rho)$ and regular $F_{l}(\chi,\rho)$ Coulomb functions with a given angular momentum $l$. It turns out that the modulus of this function squared is proportional to the proton half life [1]. For neutron emission the Coulomb parameter vanishes, but the other parameter, defining the outgoing ”neutral” spherical Hankel function, is given by the reduced radius $\rho$. The aim of this letter is to propose a similar systematics of spontaneous neutron half lives, but in terms of this quantity.
@@ 2 Theoretical background
The theory of neutron emission is an old issue, but we will remind its main details in order to fix the main concepts. The standard way to analyze a single particle (proton or neutron) unbound state is to consider it as a resonant state in continuum. This is an eigenstate of some mean field potential with complex energy and outgoing asymptotics described by the Coulomb-Hankel function [1]. For outgoing narrow resonances, called Gamow states, the real part of the energy is positive and it is called Q-value. The decay width $\Gamma$ is twice the complex part of the energy and for narrow resonances it satisfies the condition $\Gamma<<Q$. Thus, the wave function describing the emission process is given by the following ansatz
@@
defining the S-matrix $S_{l}(E)=\exp[2i\delta_{l}(E)]$, the phase shift for some narrow resonant state sharply passes through the value $\delta_{l}=\pi/2$ where only the irregular component $G_{l}$ remains. Thus, for a narrow resonance, where one has $|G_{l}(\chi,\rho)|>>|F_{l}(\chi,\rho)|$ inside the barrier [1], both complex Gamow and real scattering state descriptions give close results.
The low-lying proton eigenstates in continuum have very small decay widths and therefore are long living due to the large Coulomb potential. In this case the equivalent expression of the particle decay width of the process can be expressed by using the continuity equation as follows [1]
$\Gamma=\hbar v\frac{s_{lj}}{\Omega_{j}}|N_{lj}|^{2}~{},$ (2.3)
where $l,j$ denotes the angular momentum and total spin of the emitted particle, $s_{lj}$ the spectroscopic factor, $\Omega_{j}=j+\frac{1}{2}$ the level occupancy and $v=\sqrt{2Q/\mu}$ the particle velocity with the reduced mass $\mu=mm_{D}/m_{P}$, written in terms of the particle, daughter and parent masses. For superfluid nuclei the spectroscopic factor within the Bardeen-Cooper-Schieffer (BCS) approximation is given by the probability of the particle component in the quasiparticle state at the Fermi level, where the single particle energy lies close to the Lagrange multiplier $\epsilon_{lj}\approx\lambda$, thus giving $s_{lj}=u^{2}_{lj}=\frac{1}{2}[1+(\epsilon_{lj}-\lambda)/E_{{lj}}]\approx$ 0.5. The quantity $N_{lj}$ is called scattering amplitude and it is given by the ratio
$N_{lj}=\frac{f_{lj}(R)}{H^{(+)}_{l}(\chi,\rho)/\rho}~{},$ (2.4)
between the internal wave function normalized to unity inside nucleus $f_{lj}(R)$ and external outgoing wave, given by the Coulomb-Hankel spherical wave $H^{(+)}_{l}(\chi,\rho)/\rho$, depending upon the Coulomb-Sommerfeld parameter $\chi=2Ze^{2}/(\hbar v)$. Here, Z is the daughter charge, and nuclear reduced radius $\rho=\kappa R$ defined in terms of the linear momentum $\kappa=\mu v/\hbar$. Let us mention that the above ratio obviously does not depend upon the radius $R$ in the region where the nuclear interaction vanishes, because both internal and external functions satify the same Schrödinger equation. Let us mention that the same result is obtained within the real scattering approach [28].
At the same time, some low-lying neutron resonant states have relative large centrifugal barriers and therefore are also relative long living. Therefore the decay width can be expressed by (2.3), but the external outgoing wave for chargeless particles with $\chi=0$ is given by the spherical Hankel function $H^{(+)}_{l}(0,\rho)/\rho\equiv h^{(+)}_{l}(\rho)$ and describes the neutron motion at large distances, where the nuclear interaction vanishes.
Figure 1: The values of the centrifugal factor $|C_{l}(\rho)|^{2}$, given by Eq. (2), entering the half life (2.9) versus reduced radius $\rho$ for $l$=1 (solid line), 2 (dashed line), 3 (dot-dashed line).
The spherical Hankel function is defined in terms of spherical Neumann and Bessel functions
@@
It turn out that it can be factorized
$h^{(\pm)}_{l}(\rho)$ $=$ $h^{(\pm)}_{0}(\rho)C_{l}(\rho)~{},$ (2.6)
into monopole ($l=0$) and centrifugal ($l>0$) components
$h^{(\pm)}_{0}(\rho)$ $=$ $\frac{e^{\pm i\rho}}{\rho}$
$C_{l}(\rho)$ $=$ $\sum_{m=0}^{l}\frac{i^{m-l}}{m!(2\rho)^{m}}\frac{(l+m)!}{(l-m)!}~{}.$ (2.7)
In Fig. 1 we plotted the centrifugal factor $|C_{l}(\rho)|^{2}$ for $l$=1 (solid line), $l$=2 (dashed line) and $l$=3 (dot-dashed line) versus the reduced radius $\rho$ estimated at the geometrical nuclear radius $R_{0}=1.2A^{1/3}$
$\rho=\kappa R_{0}=\frac{(2\mu Q)^{1/2}}{\hbar}1.2A^{1/3}\approx 0.263~{}Q^{1/2}A^{1/3}~{},$ (2.8)
written in terms of the neutron Q-value and mass-number $A$. From this figure one sees that the centrifugal effect gives an effect up to two orders of magnitude per unit of angular momentum. Notice that by considering for experimental values $Q<$ 2 MeV, $A<$ 30 one obtains a relative small interval of values $\rho<1$. We also make theoretical predictions by extending the analyzed interval up to $\rho$=2, corresponding to $A=200$, allowing in this way to predict the half lives of the neutron dripline.
Thus, the theoretical half life can be written as follows
$T=\frac{\hbar\ln 2}{\Gamma}=\frac{\Omega_{j}\ln 2}{s_{lj}v}\left|\frac{h^{(+)}_{l}(R)}{f_{lj}(\rho)}\right|^{2}=\frac{\Omega_{j}\ln 2}{s_{lj}v\rho^{2}}\left|\frac{C_{l}(R)}{f_{lj}(\rho)}\right|^{2}~{}.$ (2.9)
@@ 3 Data systematics
At present the amount of neutron emission experimental data from the ground state is very limited, due to the difficulties to measure half lives of these very unstable nuclei, close to the neutron dripline [29]. Moreover, the experimental errors are relative large. This is seen from Table I, where we give the available experimental data concerning quantum numbers, Q-value, half life and error of neutron emission from ground state. In the last columns are given the corresponding references. Notice that, except for the first three lines, we assigned angular momenta according to the standard spherical shell model scheme [30], due to the fact that this quantity in general is not experimentally determined. Let us mention here that the standard closure numbers N=2, 8, 20 seems to change with the increase of the asymmetry $N-Z$. As can be seen from Fig. 1 of Ref. [31], the neutron shell closure at N=20 disapears and therefore the assignement of the angular momentum may be different from the standard scheme. Thus, we considered one exception labeled by an asterisk in Table I for $l=1$, corresponding to a much better fit to the theoretical curve, describing the angular momentum dependence of the half life.
TABLE I. Neutron experimental emission data: charge, neutron, mass numbers, angular momentum, Q-value, reduced radius estimated at the nuclear geometrical radius, half life, error and reference. $Z$ $N$ $A$ $l$ $Q$ (keV) $\rho$ $T$ (s) $\delta T$ (s) Ref. 1 3 4 1 1600 0.528 8.42(-23) - [33] 2 3 5 1 735 0.386 7.89(-22) - [34] 2 5 7 1 409 0.322 3.04(-21) 0.41(-21) [34] 4 9 13 2 510 0.442 2.70(-21) 1.80(-21) [35] 4 11 15 2 1800 0.870 7.87(-22) 2.71(-22) [36] 5 11 16 1 82 0.190 $>$4.56(-21) - [37] 5 15 20 2 1560 0.892 9.12(-22) - [38] 8 17 25 2 725 0.655 2.28(-22) ${}^{+6.84(-22)}_{-2.28(-22)}$ [39] 9 19 28 2 220 0.375 4.56(-20) - [40] 1 219 0.373 2.53(-21) ${}^{+0.72(-21)}_{-0.46(-21)}$ [41]
@@
Figure 2: Experimental half lives in logarithmic scale (in $10^{-22}$ units) versus reduced radius $\rho$ for $l$=1 (open circles) and $l=2$ (dark circles). The fitting curves are given by $T_{l}=0.1~{}|h^{(+)}_{l}(\rho)|^{2}=0.1~{}|C_{l}(\rho)|^{2}/\rho^{2}$ for angular momenta l=1,2. We also added the curve corresponding to $l=0$ and plotted predicted half lives corresponding to the interval $\rho\in[1,2]$.
These nuclei belong to the neutron dripline. They are very unstable, the largest half life being only by two orders of magnitude above the characteristic nuclear time $10^{-22}$ s. Let us stress on the fact that the amount of experimental data is very rich if one considers neutron emission from excited states populated by beta decay, defining the so-called beta delayed neutron emission [32]. The analysis of these data will be performed in a forthcoming paper.
@@
In Fig. 2 we plotted the experimental half lives in logarithmic scale (in $10^{-22}$ units) together with the simplest theoretical prediction given by (2.9), fitting these data by using one common parameter for all angular momenta
$T_{l}=~{}a|h^{(+)}_{l}(\rho)|^{2}=~{}a\frac{|C_{l}(\rho)|^{2}}{\rho^{2}}~{},$ (3.1)
where the fit parameter $a=0.1$ and the rms error $\sigma=0.5$. Thus, we neglected the influence of the factor multiplying $|h^{(+)}_{l}(\rho)|^{2}$ in Eq. (2.9). We notice a reasonable agreement between this prediction and experimental data within the displayed experimental errors. We also plotted predicted half lives corresponding to the interval $\rho\in[1,2]$.
@@
Figure 3: Experimental half lives in logarithmic scale (in $10^{-22}$ units) versus reduced radius $\rho$ divided by the centrifugal factor $|C_{l}(\rho)|^{2}$. The fitting curve represents the same ansatz as in Fig. 2 for the monopole case $T_{0}=0.1~{}|h^{(+)}_{0}(\rho)|^{2}=0.1/\rho^{2}$. We plotted predicted half-lives corresponding to the interval $\rho\in[1,2]$.
@@
Let us mention that in Ref. [10] the influence of the centrifugal barrier for proton emitters was excluded by dividing the experimental half life to the centrifugal factor $|C_{l}(\rho)|^{2}$. Here we will use the same centrifugal factor by defining the reduced half life
$T_{red}=\frac{T_{exp}}{|C_{l}(\rho)|^{2}}~{}.$ (3.2)
@@
It actually correspons to a dependence of the reduced half life upon the monopole spherical Hankel function squared. In Fig. 3 we plotted the dependence of the reduced half life in logarithmic scale (in $10^{-22}$ units) versus the reduced radius $\rho$, together with the fitting curve
$T_{0}=0.1~{}|h^{(+)}_{0}(\rho)|^{2}=\frac{0.1}{\rho^{2}}~{}.$ (3.3)
We obtained a reasonable correlation between this curve and experimental points.
@@
Figure 4: Experimental formation probability given by the ratio $T_{exp}/T_{l}$, where $T_{l}$ is gicen by (3.1), in logarithmic scale.
@@
Let us here mention that the role of the particle emission formation probability, prportional to the wave function squared $|f_{lj}(R)|^{2}$, was extensivelly investigated. In the recent review [13] various versions of the shell model calculations in this region were analyzed, starting with the Kuo-Brown code [42]. The particle emission processes were analyzed by using the resonant states in a deformed Woods-Saxon plus Coulomb mean field [43, 44, 45], as well as more sophisticated open quantum system formalism for the description of weakly bound nuclei far from the valley of stability [46, 47].
@@
In order to evidence the role of the neutron wave function we plotted in Fig. 4 the so-called ”experimental formation probability”, given by the ratio between the experimental half life and the best fit of the centrifugal penetration factor $f_{exp}(\rho)=T_{exp}/[0.1/\rho^{2}]$. This quantity is obviously proportional to the neutron wave function squared $|f_{lj}(R)|^{2}$ entering the definition of the half life ((2.9). The result in Fig. 4 is nothing else that the ratio between experimental points and the corresponding values of the curve in Fig. 3. Notice that the rms error is $\sigma=0.5$, corresponding to a factor about three in the half life. Therefore the role of the neutron wave function is not so important for a rough estimate of the neutron emission half life for transitions connecting ground states.
@@
Figure 5: Logarithm of the half life versus energy of l=2 (upper curve) and l=1 (lower curve) resonant levels in a spherical Woods-Saxion mean field for 28F. The vertical line represents the experimental Q-value and the two horizontal lines the experimental half lives given by the last two lines of Table I.
@@
On the other hand, it turns out that the most sensitive ingredient is the angular momentum of the emitted neutron, similar to the proton emission systematics [10]. This is shown in Fig. 2 and also in Fig. 5, where we plotted the theoretical half life of a resonant state above the Fermi surface with angular momentum l=1 (lower curve) and l=2 (lower curve) versus its energy for 28F. Notice here two orders of magnitude difference in half lives between these curves at the experimental Q-value. These states were generated by using a spherical Woods-Saxon meand field [1] with universal parametrisation [48, 49] for 28F. We used the theoretical decay width given by (2.3) and changed the energy of the resonant state by using the real part of the central potential. It turns out that this relation is not sensitive to the change of other parameters of the used nuclear mean field. This feature is also general for proton emission processes connecting ground states [11]. By vertical dotted line we denoted the experimental Q-value and by horizontal dotted lines the two experimental values of the half life mentioned in the last two lines of Table I. Let us mention that our theoretical prediction for l=1 is much closer to the experintal value, thus confirming the above statement aboout the island of inversion in this region and magicity of 28O.